OpenStudy (anonymous):

Sam has a total of 58 DVD's and CD's. If the number of CD's is two more than three times the number of DVD's, how many CD's does he have? a. 42 b. 14 c. 44 d. 12

OpenStudy (anonymous):

again its the same...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

try to think xD

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let D be the number of DVDs. Then C, the number of CDs is C = 3D + 2 The total is 58, so C + D = 58 3D + 2 + D = 58 4D = 56 D = 14 C = 3(14) + 2 = 42 + 2 = 44 So he has 14 DVDs and 44 CDs

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

yw